An elevator car whose floor to ceiling distance is equal to 2.7m starts ascending with constant acceleration of 1.2m/s . Two sec after it starts; a bolt begins to fall from the ceiling of the elevator. Find 
(a) the bolts free fall time; 


(b) The displacement and the distance covered by the bolt during the fall w.r.t. ground. 

Grade:

2 Answers

md salik
11 years ago

s=ut+1/2 at2

2.7=1/2(g+a)*t2

2.7=11/2*t2

11 t2=5.4

approx-

t2=0.5

t=2.2 sec

Satyam Mahato
11 years ago

part 1 is correct.

But what is the answer of 2nd part.